plane leaves an airport traveling at an average speed of 240 kilometers per hour. how long will it take a second plane traveling the same route at an average speed of 600 kilometers per hourto catch up with the first plane if it leaves 3 hours later?
Answer by checkley71(8403)

240T=600(T-3)
240T=600T-1800
240T-600T=-1800
360T=-1800
T=1800/360
T=5 HOURS THE FASTER PLANE WILL CATCH UP WITH THE SLOWER PLANE.
PROOF
240*5=600*2
1200=1200